The University Letter Style:

           csletter and euletter

Irene Orr

(Updated by Martin Reddy)

January 23, 1997

1 Introduction

This is a modification of the standard L^{A}TEX “letter” style for the University of Edinburgh’s corporate visual identity. It is customised for the Department of Computer Science, to take into account the departmental information, the method of including the logo, and the header font selection.

It conforms to the generic style for Edinburgh University letters, following the visual identity style guidelines. An appropriate departmental style file defines the addresses, professors, etc.

2 Usage

The csletter class is invoked in the same way as any other class; that is, you begin your document with \documentclass{csletter}. Within the actual document, you should include all of your letter text within a letter environment, i.e. \begin{letter}{<address>} ...

\end{letter}.

Do NOT use the \address macro to set the address; it will be ignored. There are several macros for user-level customisation. These are:

     \email{address}      Sets the user's email address
     \extn{number}        Sets the new phone extension
     \telex{number}       Sets the telex address
     \href{url}           Sets a World Wide Web page URL

If any of these are given an empty argument, the corresponding address line will be omitted.

There is also the macro \code{number} which sets the phone prefix (useful for international style). The macro \crestversion{num} selects the crest version:

      0  -  normal monochrome
      1  -  no lines
      2  -  reversed
      3  -  2 colour
      4  -  full colour

The commands \ourref{num} and \yourref{num} can be used in the recipient’s address to set up reference numbers for the letter.

The command \newcc{\name}{intro:} creates a new \cc-like command, with the text ”intro:” at the start.

3 Class Options

headedpaper : will omit most of the address, and just fill in the gaps for pre-printed letter paper.

onepage : will alter the margins to fit as much as possible onto the first page. This is useful for letters in which the signature would otherwise go on the second page on its own.

cm : will use Computer Modern for the appropriate parts of the address (perpetua is the default).

4 Customisation

The general University style is defined by the class euletter. The class csletter is built on top of euletter and defines the departmental specific information such as the address, telephone/fax numbers, professors, etc.

Other departments, or users wishing to customise the letter style for personal use, should leave the euletter class as-is and update only the contents of csletter. For individual users, this can be done by copying the csletter file into their $(HOME)/tex/latex directory and modifying this file appropriately (N.B. in the Department of Computer Science, csletter is merely a wrapper for a class called csletr. It is therefore this latter file that should be updated by local users).